((y+7)/y) = (8/7)
is this even possible?

Answer by dfrazzetto(283)

((y+7)/y) = (8/7)
Cross multiply:
8y = 7(y+7)
8y = 7y + 49
8y - 7y = 49
y = 49
CHECK:
(49+7) / 49 = 56/49 = 8/7
It is possible ;-)